# GroundGazer: Camera-based indoor localization of mobile robots with millimeter accuracy at low cost

### v3

@@ Abstract
Highly accurate indoor localization systems with mm positioning accuracy are currently very expensive. They include range finders (such as LiDAR), tachymeters, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost, easy to set up, portable, robust, scalable to large areas and robot swarms, and potentially extendable to 3D position and orientation estimation.
@@ I Introduction
High-accuracy planar indoor localization applications, e.g. docking in warehousing and benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy, require localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading estimation accuracy with high robustness against occlusion of the position references on the ground at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
Before describing the GG system, we give an introduction into currently existing indoor localization methods and sensor technologies for AMRs with mm positioning accuracy. For an overview of general indoor localization technologies, we refer to the surveys in [review_indoor_pos_habil] and [review_indoor_loc]. Comprehensive overviews of optical indoor localization systems can be found in [Survey_Optical_pos] and [cv_loc_survey].
@@ I-A Localization technologies with mm positioning accuracy
Motion capture systems consist of passive infrared (IR) markers placed on the to be tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. Motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/. are very popular in robotics and industry. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require line of sight (LOS), multiple high-end IR cameras, and become very expensive if large areas are to be observed. Further, the IR cameras not installed around the AMR complicate the time synchronization when the motion capture system is used as reference for another localization system with sensors installed on the AMR. Highly accurate time synchronization is required, since otherwise the mm localization accuracy is lost through badly matched time stamps.
@@
A different, less frequently used solution for 3D position estimation with mm or sub-mm accuracy are tachymeters, e.g. used as 3D position reference system in [esparagos]. Tachymeters enable long range sensing, but like motion capture systems, they are expensive and require LOS and more complex time synchronization mechanisms. Another, novel positioning system with long distance capabilities based on interferometry is developed by Koherent [^2] Available at: https://koherent.io/.
@@
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [hawkeye] or triangulation [mm_siemens] for positioning.
@@
In acoustic ultrasound systems, small timing errors in the measured time of flight (ToF) of the signals or in the synchronization between anchors cause lower positioning errors than in RF based ultra wideband (UWB) systems. The reason lies in the much lower propagation speed of acoustic waves. The company ZeroKey claims real-time and up to 1.5 mm 3D positioning accuracy with their commercial Quantum RLTS 2.0 system, given 6 acoustic anchors with ideal geometric placement [zerokey]. Positioning is performed with multiple active acoustic anchors and ToF measurements between anchors and the to be tracked AMR and by applying MLAT for localization. Since no product prices are available on their website or data sheets, we believe it likely not being a low-cost solution for AMR localization.
One of the most widely used localization sensor is LiDAR [lidar]. While high-end LiDAR systems can achieve 3D mm accuracy, they are also expensive and need precise maps of the environment through costly measurement campaigns to achieve accurate, absolute position estimates. Further, their accuracy degrades with dynamic changes of the environment that are not contained in the previously recorded reference map.
@@
Deep learning based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system has recently also shown mm accuracy in a small area and under LOS conditions [csi]. However, similar as with LiDAR, the performance of such system degrades with changes in the environment, causing domain shifts between the measured fingerprints during inference and the pre-recorded fingerprint database.
@@
Fiducial markers pose another solution for highly precise localization. In [ODOP], the combination of high-end binocular cameras and custom markers achieves sub-mm localization accuracy. Downsides are again the high cost of the cameras and the required LOS. The authors of [QR_docking_submm] proposed a docking system with sub-mm 2D positioning accuracy using quick response (QR) codes and an RGB camera. Although low-cost, the very short camera-QR code distance and the need for many QR codes limit its use for general AMR localization.
@@
Additional systems based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) with (sub-)mm accuracy can be found in [Survey_Optical_pos] and [mautz_alternative_systems].
@@ I-B Contributions
We introduce a novel mm accuracy localization approach with the intended application to serve as reference for less accurate indoor positioning systems. Our GG prototype described in this paper requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost off-the-shelve components. Currently, it only allows for planar position ($x$, $y$) and heading/yaw $\theta$ estimation, which is sufficient for our application as reference system for planar AMR position and heading estimation with known or irrelevant height, roll and pitch. In Sec. VI, we discuss an adaption that yields 3D position and orientation estimates, but has higher computational complexity and possibly lower accuracy.
The idea for our GG system originates from [ground_truth_laser]. They developed a fiducial marker localization system that requires a chessboard floor in addition to the fiducial markers, where a car is equipped with multiple fiducial markers placed on a plate mounted at the front of the car. To evaluate the positioning accuracy of this system, they installed two point laser diodes that project two points from the front and the back of the car onto the chessboard floor beneath it. The corner coordinates of the chessboard serve as position references and they measure the two laser point locations on the floor as ground truth by hand. The laser point at the front of the car is used as vehicle reference position and the heading is derived from the orientation of the line connecting both points w.r.t. to the chessboard. Our GG system builds on that idea, resulting in a highly accurate, camera-based indoor localization system.
@@ I-C System overview
((b)) 2D position and heading estimation using the detected crosshair (P5) and chessboard square corners (P1-P4). The globally known square corner coordinates through the QR codes enable accurate, global positioning.
@@
The localization system works by detecting the crosshair in a square of the chessboard. We specifically selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The laser diode is installed such that the AMR moves in the direction of the vertical crosshair line, thus the angle between the vertical crosshair line and the chessboard lines in $x$ direction gives the AMR’s heading. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
@@
Since the chessboard corner coordinates and the crosshair orientation are ambiguous to the camera which only observes a small, local area of the periodic chessboard pattern, we additionally equip the floor with different QR codes as global position and orientation references. All QR codes are oriented in the same direction, parallel to the chessboard lines, and their corners can therefore be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the distance in $x$- and $y$-direction between a detected QR code square center and the center of the square in which the crosshair lies (neglecting the trivial case where they are identical), we can further identify the crosshair square and its known corner coordinates without requiring a QR code in each square. Measuring the relative position of the crosshair w.r.t. its square corners (at least three corners are required to build a square coordinate system, but we use four corners as known from a detected square) gives the AMR’s position and the orientation of the vertical crosshair line w.r.t. the chessboard lines in $x$ direction its heading. The AMR position and heading is set as the crosshair position and orientation. An image describing the AMR positioning is shown in Fig. 1(b).
@@
High position and heading accuracy is enabled by placing the camera close to the ground and by precise estimation of the crosshair center and the crosshair lines, and the square corners and the chessboard lines. For this prototype, a camera with fisheye lens was chosen because of their large FoV, allowing observation of multiple chessboard squares with a small camera-ground distance. The whole processing chain, which transforms the initial fisheye image, detects the QR codes, the crosshair, and the chessboard squares, and finally estimates and smooths the AMR’s position and heading, is described in Sec. III.
II Related works with ground facing cameras
@@
In this section, we focus on closely related systems that employ ground facing cameras for global localization. A general overview of indoor positioning systems with mm accuracy for AMRs was given in Sec. I-A.
@@
One approach to global localization using ground facing cameras is by avoiding external infrastructure, i.e. localization using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) [prob_robotics], but we are interested in absolute, global positioning. Such positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [high_precision_matching, streetmap, binary_matching, feature_matching_newest]. The advantage of this is the applicability in (ideally) any kind of environment without requiring new infrastructure, making it a practical method e.g. for warehousing. However, feature-matching requires an expensive measurement campaign for collection of reference images of the floor with accurate ground truth, e.g. from LiDAR or motion capture systems. These approaches have also shown varying reliability for different surfaces [high_precision_matching] and the position errors are in the low cm range (ignoring outliers, but also not yet using temporal correlations) and the heading errors in the low degree range [feature_matching_newest]. Position errors further seem to increase with coverage area (map size) [feature_matching_newest]. The commercial feature-matching based systems by Accerion are reported to deliver sub-mm accuracy, but only for repeated tasks such as docking, where the task was accurately executed before, while stating cm accuracy for global localization [^3] Due to the limited product specifications on the company website (no published data sheets), we are relying on the information given on their Youtube channel found at https://www.youtube.com/@accerion4564.. While feature-matching approaches require no additional infrastructure, we argue that using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements.
@@
Various approaches have considered placing infrastructure on the floor for localization. In [floor_aruco], ArUco markers [ArUco] are distributed on the floor and combined with odometry in an Extended Kalman filter (EKF) for localization. However, they don’t give localization accuracy results. QR codes are also commonly applied, e.g. for AMRs in [QR_amr]. In [QR_drones], a drone with a ground facing camera was localized with a QR code grid on the floor and using the nearest QR code for position and heading estimation, showing cm accuracy. Different fiducial marker types are benchmarked regarding their detection behavior in ground based localization in [fiducial_floor_det].
@@
It should be noted, however, that we believe mm accuracy with floor installations of ArUco, QR codes, or similar fiducial markers is possible. With ceiling installations of ArUco markers and much larger camera-marker distances, we could already achieve low cm accuracy in previous work [icar25_fiducial] and [QR_docking_submm] reports sub-mm accuracy with very small camera-QR code distance. Similar commercial systems can also be found. One example is the company Hikrobot, which provides AMRs and vision based localization solutions using a visual marker grid on the floor for warehousing, logistics etc. [^4] Available at: https://www.hikrobotics.com/en/.
@@
Opposed to existing systems, we can deliver absolute 2D position and heading estimates with mm and sub-degree accuracy at low cost and with minimal setup. Our processing can likely also be sped-up through further optimization as discussed in Sec. VI to enable near real-time mm accuracy during online processing with common AMR velocities of $1\text{\,}\frac{m}{s}$, factoring in additional delays from exposure time and image transmission to the processor. This online bottleneck might be solved by adding high-frequency odometry. Neither SLAM nor feature-matching can produce this due to higher processing times and delayed position estimates. To the best of our knowledge, no system like GG exists right now for indoor AMR localization that offers a similar combination of high accuracy, low cost, simplicity, easy set up, scalability, portability, and fast processing.
@@ III-A Image transformation
The first step of our processing is the transformation of the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image needs to be undistorted. We apply the built-in undistort function from OpenCV [opencv_library] [^5] If not specified otherwise, we always refer to OpenCV functions for the remainder of the paper.. This function takes as input the distorted fisheye image, the intrinsic matrix $\mathbf{K}$ of the camera, and its distortion coefficient vector $\underline{d}$, and outputs the undistorted image. The intrinsic matrix $\mathbf{K}$ and the distortion coefficient vector $\underline{d}$ have to be determined beforehand by camera calibration. We used a printed chessboard pattern and the calibrateCamera function for this task. A distorted fisheye image and the result after undistortion are shown in Fig. 2(a) and Fig. 2(b), respectively. The second step transforms the undistorted image to a top-down view. This can be realized with the warpPerspective function, which takes the undistorted image and the homography matrix as input and gives the top-down view. The homography relates the undistorted and the top-down view. It can be estimated with the findHomography function and corresponding image points of the two views. We take four known corner points of the most central chessboard square with a size of (300, 300) pixels after transformation for the wanted top-down view and select four corresponding chessboard corner points from the undistorted image. The transformed image from the top-down perspective is given in Fig. 2(c).
@@
Figure 2: Transformation of the original image in a) through undistortion in b) and perspective transformation in c).

### v4

@@ Abstract
Highly accurate indoor localization systems with mm positioning accuracy are currently very expensive. They include laser trackers, total stations, and motion capture systems relying on multiple high-end cameras. In this work, we introduce a high-accuracy, planar indoor localization system named GroundGazer (GG) for autonomous mobile robots (AMRs). GG estimates the AMR’s position with mm and its heading with sub-degree accuracy. The system requires only a monocular (fisheye) camera, a chessboard floor, and an optional laser diode. Our system is simple and low-cost, easy to set up, portable, robust, scalable to large areas and robot swarms, and extendable to 3D position and orientation estimation.
@@ I Introduction
Benchmarking lower-accuracy indoor positioning systems with up to cm localization accuracy requires localization approaches with reliable mm positioning accuracy. Currently, such systems are very costly (generally $1000\,\mathdollar$, with costs rising quickly with increasing size of the coverage area). In this work, we introduce a novel camera-based indoor localization system for AMRs which achieves mm localization and sub-degree heading accuracy at multitudes lower cost than existing systems of similar precision. We name it GroundGazer, as the camera placed on the AMR gazes at the passive position references on the ground.
Before describing the GG system, we give an introduction into currently existing and researched indoor localization technologies for AMRs with mm positioning accuracy. For an overview of general indoor localization technologies, we refer to the surveys in [24] and [34]. Comprehensive overviews of optical indoor localization systems can be found in [22] and [25].
@@ I-A Localization technologies with mm positioning accuracy
Motion capture systems consist of passive infrared (IR) markers placed on the to be tracked AMR. With one marker, 3D positioning is possible. A minimum of three non-collinear markers enable 3D position and orientation estimation. Motion capture systems like the commercial systems by Qualisys and Vicon [^1] Available at: https://www.qualisys.com/ and https://www.vicon.com/. are very popular in robotics and industry. While capable of sub-mm 3D localization and sub-degree orientation estimation, they require line of sight (LOS), multiple high-end IR cameras, and become very expensive if large areas are to be observed. Further, the IR cameras not installed around the AMR complicate the time synchronization when the motion capture system is used as reference for another localization system with sensors installed on the AMR. Highly accurate time synchronization is required, since otherwise the mm accuracy is lost through badly matched time stamps.
@@
Total stations (tachymeters) pose a another solution for 3D position estimation with mm or sub-mm accuracy, e.g. used as 3D position reference system in [11]. Total stations enable long range sensing, but like motion capture systems, they are expensive and require LOS and more complex time synchronization mechanisms.
A novel sub-mm positioning system with long distance capabilities based on radio signals and interferometry that has recently seen real-world deployments is developed by Koherent [^2] Available at: https://koherent.io/.
@@
Radio frequency (RF)-based backscatter systems using frequency modulated continuous wave (FMCW) radar and (semi-)active reflective tags (e.g. Van-Atta arrays) that modulate the reflected radar signal can also achieve 3D mm localization accuracy using high-precision range estimates of the tags and multilateration (MLAT) [1] or triangulation [33] for positioning. The system in [1] could also demonstrate good performance with large radar-tag distance.
@@
In acoustic ultrasound systems, small timing errors in the measured time of flight (ToF) of the signals or in the synchronization between anchors cause lower positioning errors than in RF-based ultra wideband (UWB) systems. The reason lies in the much lower propagation speed of acoustic waves. The company ZeroKey claims real-time and up to 1.5 mm 3D positioning accuracy with their commercial Quantum RLTS 2.0 system, given 6 acoustic anchors with ideal geometric placement [35]. Positioning is performed with multiple active acoustic anchors and ToF measurements between anchors and the to be tracked AMR and by applying MLAT for localization.
While ToF-based localization with UWB systems only gives dm to cm accuracy, recent studies show mm accuracy by combining initial, coarse ToF-based range estimates with refined, ambiguity-resolved phased-based range estimates [21].
@@
Deep learning-based channel state information (CSI) fingerprinting with a massive multiple input multiple output (MIMO) system could recently achieve mm accuracy in a small area and under LOS conditions [20]. However, the performance of such system degrades with changes in the environment, causing domain shifts between the measured fingerprints during inference and the pre-recorded fingerprint database.
@@
Fiducial markers pose another solution for highly precise localization. In [14], the combination of high-end binocular cameras and custom markers achieves sub-mm localization accuracy. Downsides are again the high cost of the cameras and the required LOS. The authors of [4] proposed a docking system with sub-mm 2D positioning accuracy using quick response (QR) codes and an RGB camera. Although low-cost, the very small camera-QR code distance and the need for many QR codes limit its use for general AMR localization.
@@
Other systems with (sub-)$m$ accuracy based on visual markers, laser projections, magnetic fields, laser trackers, pseudolites, and infrared laser beams (iGPS) can be found in [22] and [23].
@@ I-B Contributions
We introduce a novel mm accuracy localization approach with the intended application to serve as reference for less accurate indoor positioning systems. The GG prototype described in this paper requires only a chessboard floor (or similar grid pattern), a camera (here a global shutter RGB fisheye camera), and an optional laser diode, i.e. low-cost off-the-shelve components. Currently, it only allows for planar position ($x$, $y$) and heading $\theta$ estimation, which is sufficient for our application as reference system for planar AMR position and heading estimation with known or irrelevant height, roll and pitch. In Sec. VI, we discuss an adaption that yields 3D position and orientation estimates, but has higher computational complexity and possibly lower accuracy.
The idea for our GG system originates from [3]. They developed a fiducial marker localization system that requires a chessboard floor in addition to the fiducial markers, where a car is equipped with multiple fiducial markers placed on a plate mounted at the front of the car. To evaluate the positioning accuracy of this system, they installed two point laser diodes that project two points from the front and the back of the car onto the chessboard floor beneath it. The corner coordinates of the chessboard serve as position references and they measure the two laser point locations on the floor as ground truth by hand. The laser point at the front of the car is used as vehicle reference position and the heading is derived from the orientation of the line connecting both points w.r.t. to the chessboard. Our GG system builds on that idea, resulting in a highly accurate, camera-based indoor localization system.
@@ I-C System overview
((b)) 2D position and heading estimation using the detected crosshair (P5), the chessboard square corners (P1-P4), and the $x$-direction chessboard lines. The globally known square corner coordinates through the QR codes enable accurate, global positioning.
@@
The localization system works by detecting the crosshair in a square of the chessboard. We specifically selected a blue and white chessboard floor and a green laser diode, as green lasers are most easily visible and other commercially available black and white chessboard floors complicate the crosshair detection, caused by the large difference in laser reflective behavior between black and white squares. The camera image is calibrated such that the heading of the AMR can be estimated based on the orientation of the chessboard in the image. The corner coordinates of the chessboard squares are known and serve as passive position references for AMR localization.
@@
Since the chessboard corner coordinates are ambiguous to the camera which only observes a small, local area of the periodic chessboard pattern, we additionally equip the floor with different QR codes as global position and orientation references. All QR codes are oriented in the same direction, parallel to the chessboard lines, and their corners can therefore be used to extract the $x$- and $y$-directions of the chessboard lines in global coordinates. The QR code IDs allow for identification of the QR code squares and global positioning. By measuring the distance in $x$- and $y$-direction between a detected QR code square center and the center of the square in which the crosshair lies (neglecting the trivial case where they are identical), we can further identify the crosshair square and its known corner coordinates without requiring a QR code in each square. Measuring the relative position of the crosshair w.r.t. its square corners (at least three corners are required to build a coordinate system, but we use four corners as known from a detected square) gives the AMR’s position and the angle between the chessboard lines in $x$-direction and the vertical image direction gives its heading. The AMR position is assumed as the crosshair position. An image describing the AMR positioning is shown in Fig. 1(b).
@@
High position and heading accuracy is enabled by placing the camera close to the ground and by precise estimation of the chessboard square corners, the chessboard lines, and the crosshair center. For this prototype, a camera with fisheye lens was chosen because of their large FoV, allowing observation of multiple chessboard squares with a small camera-ground distance. The whole processing chain, which transforms the initial fisheye image, detects the QR codes, the crosshair, and the chessboard square corners, and finally estimates the AMR’s position and heading, is described in Sec. III.
II Related works with ground gazing cameras
@@
In this section, we focus on closely related systems that employ ground gazing cameras for global localization. A general overview of indoor positioning systems with (sub)-mm accuracy for AMRs was given in Sec. I-A.
@@
One approach to global localization using ground gazing cameras is by avoiding external infrastructure, i.e. localization using only the given ground texture. Relative positioning can be realized through simultaneous localization and mapping (SLAM) [32], but we are interested in absolute, global positioning. Such positioning is possible with feature-matching of images collected during inference with a pre-collected image database of the ground [17, 36, 7, 12, 6]. The advantage of this is the applicability in (ideally) any kind of environment without requiring new infrastructure, making it a practical method e.g. for warehousing. However, feature-matching requires an expensive measurement campaign for collection of reference images of the floor with accurate ground truth, e.g. from motion capture systems. These approaches have also shown varying reliability for different surfaces [36] and the position errors are in the low cm range (ignoring outliers, but also not yet using temporal correlations) and the heading errors are in the low degree range [6]. Positioning errors further seem to increase with coverage area (map size) [6]. The commercial feature-matching-based systems by Accerion are reported to deliver sub-mm accuracy, but only for repeated tasks such as docking, where the task was accurately executed before, while stating cm accuracy for global localization [^3] Due to the limited product specifications on the company website (no published data sheets), we are relying on the information given on their Youtube channel found at https://www.youtube.com/@accerion4564.. For relative localization, features of successive images can similarly be matched to build a vision-based odometer [27]. While feature-matching approaches require no additional infrastructure, we argue that using known position references is the more suitable approach for reference systems due to higher accuracy, robustness, easier setup, and lower hardware requirements.
@@
Various approaches have considered placing infrastructure on the floor for absolute, global localization. In [37], ArUco markers are distributed on the floor and combined with odometry in an Extended Kalman filter (EKF) for AMR localization. However, they don’t give localization accuracy results. AMR localization with a simulated grid of ApriTags yields cm accuracy in [18]. QR codes are also commonly applied, e.g. for a heeled robotic walking assistant in [26], achieving dm accuracy with additional odometry and gyroscope data. In [16], a drone with a ground facing camera is localized with a QR code grid on the floor and using the nearest QR code for position and heading estimation, showing cm accuracy. Different fiducial marker types are benchmarked regarding their detection behavior in ground-based localization in [8].
@@
It should be noted, however, that we believe low-cost mm accuracy with floor installations of ArUco, AprilTag, QR codes, or similar fiducial markers is possible. With ceiling installations of ArUco markers and much larger camera-marker distances, we could achieve low cm accuracy in previous work [15] and [4] reports sub-mm accuracy with very small camera-QR code distance. Similar commercial systems can also be found. One example is the company Hikrobot, which provides AMRs and vision-based localization solutions using a visual marker grid on the floor for warehousing, logistics etc. [^4] Available at: https://www.hikrobotics.com/en/. The Hercules AMR by Amazon robotics is another commercial system using such floor marker grid for navigation.
@@
Opposed to existing systems, we can deliver absolute 2D position and heading estimates with mm and sub-degree accuracy at low cost and with minimal setup. Our processing can likely also be sped-up through further optimization as discussed in Sec. VI to enable near real-time mm accuracy during online processing with common AMR velocities of $1\text{\,}\frac{m}{s}$, factoring in additional delays from exposure time and image transmission to the processor ($10\text{\,}\mathrm{m}\mathrm{s}$ delay at $1\text{\,}\frac{m}{s}$ causes $10\text{\,}\mathrm{m}\mathrm{m}$ error, i.e. cm online accuracy). This online bottleneck might be solved by integrating a high-frequency inertial measurement unit (IMU). Neither current SLAM nor feature-matching approaches can produce this due to higher processing times. To the best of our knowledge, no system like GG exists right now for indoor localization of AMRs that offers a similar combination of high accuracy, low cost, simplicity, easy setup, scalability, portability, and fast processing.
@@ III-A Image transformation
The first step of our processing is the transformation of the distorted and skewed image collected by the fisheye camera to a top-down view of the chessboard floor. This step consists of two transformations. First, the fisheye image needs to be undistorted. We apply the built-in undistort function from OpenCV [5] [^5] If not specified otherwise, we always refer to OpenCV functions for the remainder of the paper.. This function takes as input the distorted fisheye image, the intrinsic matrix $\mathbf{K}$ of the camera, and its distortion coefficient vector $\underline{d}$, and outputs the undistorted image. The intrinsic matrix $\mathbf{K}$ and the distortion coefficient vector $\underline{d}$ have to be determined beforehand by camera calibration. We used a printed chessboard pattern and the calibrateCamera function for this task. A distorted fisheye image and the result after undistortion are shown in Fig. 2(a) and Fig. 2(b), respectively. The second step transforms the undistorted image to a top-down view. This can be realized with the warpPerspective function, which takes the undistorted image and the homography matrix as input and gives the top-down view. The homography relates the undistorted and the top-down view. It can be estimated offline with the findHomography function and corresponding image points of the two views. We define a square of size (300, 300) pixels with horizontal and vertical edges for the wanted top-down view after perspective transformation and we select the four corresponding chessboard corner points from the undistorted image. The perspective transformed top-down image is given in Fig. 2(c). The AMR orientation has to match the direction of the calibration square, i.e. in our case $\theta=90^{\circ}$, to align the heading with the transformed chessboard orientation used for heading estimation. Alternatively, one could use the angle between a similarly calibrated, vertical crosshair line that points in heading direction and the $x$-direction chessboard line. This option would make heading estimation independent of the camera orientation and thus of camera vibrations. However, it is still susceptible to lesser laser diode vibrations and it makes heading estimation dependent on the crosshair line detection, which is worse than the chessboard line detection due to broader edges, crosshair distortion on uneven ground etc., and therefore not chosen.
@@
Figure 2: Transformation of an original image recorded with an AMR heading slightly below $90^{\circ}$ in a) through undistortion in b) and perspective transformation in c).